# Inference for Multiple Change-points in Piecewise Locally Stationary Time Series

### v1

@@ Abstract
Keywords: kink; local periodogram; time-varying autoregressive model; scan statistics; structural break;.
@@ 3.1 Problem Settings
Denote the first derivative of the spectral density as $f^{(1)}_{k}(u,\lambda)=\frac{\partial{}}{\partial{u}}f_{k}(u,\lambda)$. Thus, the time-varying spectral density for the whole process is $f(u,\lambda)=\sum_{k=1}^{m+1}f_{k}(u,\lambda)\mathbbm{1}_{\{r_{k-1}<u\leq r_{k}\}}(u)$ and its first derivative is $f^{(1)}(u,\lambda)=\sum_{k=1}^{m+1}f^{(1)}_{k}(u,\lambda)\mathbbm{1}_{\{r_{k-1}<u\leq r_{k}\}}(u)$. Note that jumps and kinks in parameter curves for tvAR models correspond to jumps and kinks in the spectral density function $f(u,\lambda)$ for some frequency $\lambda\in[-\pi,\pi]$ respectively, see Section A4 in the Supplementary Materials for some illustrations. It will be shown in Section 4 that jumps and kinks lead to different asymptotic behaviors of the change-point estimator.
@@ 3.4.2 Final Adjustment and Confidence Interval Construction for a Kink
and hence $\hat{\boldsymbol{\eta}}_{k}^{(3)}=\operatorname*{argmax}_{\boldsymbol{\eta}_{k}}S_{k,T}(\boldsymbol{\eta}_{k})$. Denote the covariance of the first derivative of $S_{k,T}(\boldsymbol{\eta}_{k})$ with respect to $\boldsymbol{\eta}_{k}$, and the second derivative of $S_{k,T}(\boldsymbol{\eta}_{k})$ with respect to $\boldsymbol{\eta}_{k}$, as $\nabla_{\boldsymbol{\eta}}S_{k,T}(\boldsymbol{\eta}_{k})\nabla_{\boldsymbol{\eta}}S_{k,T}(\boldsymbol{\eta}_{k})^{\prime}$ and $\nabla_{\boldsymbol{\eta}}^{2}S_{k,T}(\boldsymbol{\eta}_{k})$, respectively; see Section A3 in Supplementary Materials for their expressions. Then, from the asymptotic property in (4.2), the asymptotic variance $\mathbf{\Sigma}_{k}$ of $\hat{\boldsymbol{\eta}}_{k}^{(3)}$ can be estimated by the plug-in estimator $\hat{\mathbf{\Sigma}}_{k,T}=\hat{\mathbf{D}}_{k,T}^{-1}\hat{\mathbf{G}}_{k,T}\hat{\mathbf{D}}_{k,T}^{-1}$, where $\hat{\mathbf{G}}_{k,T}=\nabla_{\boldsymbol{\eta}}S_{k,T}(\hat{\boldsymbol{\eta}}_{k}^{(3)})\nabla_{\boldsymbol{\eta}}S_{k,T}(\hat{\boldsymbol{\eta}}_{k}^{(3)})^{\prime}$ and $\hat{\mathbf{D}}_{k,T}=\nabla_{\boldsymbol{\eta}}^{2}S_{k,T}(\hat{\boldsymbol{\eta}}_{k}^{(3)})$ . Hence, a marginal $100\times(1-\alpha)\%$ confidence interval for $\tau_{k}^{0}$ can be constructed by $\hat{\tau}_{k}^{(3)}\pm z_{\alpha/2}\frac{T\sqrt{[\hat{\mathbf{\Sigma}}_{k,T}]_{11}}}{\sqrt{\hat{\tau}_{k+1}^{(3)}-\hat{\tau}_{k-1}^{(3)}}}$, where $z_{\alpha/2}$ is the $(1-\alpha/2)$ quantile of standard normal distribution and $\hat{\mathbf{\Sigma}}_{k,T}]_{11}$ is the element at the last row and last column of the matrix $\hat{\mathbf{\Sigma}}_{k,T}$.

### v2

@@ Abstract
Keywords: kink; local periodogram; time-varying autoregressive model; scan statistics; structural break.
@@ 3.1 Problem Settings
Denote the first derivative of the spectral density as $f^{(1)}_{k}(u,\lambda)=\frac{\partial{}}{\partial{u}}f_{k}(u,\lambda)$. Thus, the time-varying spectral density for the whole process is $f(u,\lambda)=\sum_{k=1}^{m+1}f_{k}(u,\lambda)\mathbbm{1}_{\{r_{k-1}<u\leq r_{k}\}}(u)$ and its first derivative is $f^{(1)}(u,\lambda)=\sum_{k=1}^{m+1}f^{(1)}_{k}(u,\lambda)\mathbbm{1}_{\{r_{k-1}<u\leq r_{k}\}}(u)$. Note that jumps and kinks in parameter curves for tvAR models correspond to jumps and kinks in the spectral density function $f(u,\lambda)$ for some frequency $\lambda\in[-\pi,\pi]$ respectively, see Section 8.4 in the Supplementary Materials for some illustrations. It will be shown in Section 4 that jumps and kinks lead to different asymptotic behaviors of the change-point estimator.
@@ 3.4.2 Final Adjustment and Confidence Interval Construction for a Kink
and hence $\hat{\boldsymbol{\eta}}_{k}^{(3)}=\operatorname*{argmax}_{\boldsymbol{\eta}_{k}}S_{k,T}(\boldsymbol{\eta}_{k})$. Denote the covariance of the first derivative of $S_{k,T}(\boldsymbol{\eta}_{k})$ with respect to $\boldsymbol{\eta}_{k}$, and the second derivative of $S_{k,T}(\boldsymbol{\eta}_{k})$ with respect to $\boldsymbol{\eta}_{k}$, as $\nabla_{\boldsymbol{\eta}}S_{k,T}(\boldsymbol{\eta}_{k})\nabla_{\boldsymbol{\eta}}S_{k,T}(\boldsymbol{\eta}_{k})^{\prime}$ and $\nabla_{\boldsymbol{\eta}}^{2}S_{k,T}(\boldsymbol{\eta}_{k})$, respectively; see Section 8.3 in Supplementary Materials for their expressions. Then, from the asymptotic property in (4.2), the asymptotic variance $\mathbf{\Sigma}_{k}$ of $\hat{\boldsymbol{\eta}}_{k}^{(3)}$ can be estimated by the plug-in estimator $\hat{\mathbf{\Sigma}}_{k,T}=\hat{\mathbf{D}}_{k,T}^{-1}\hat{\mathbf{G}}_{k,T}\hat{\mathbf{D}}_{k,T}^{-1}$, where $\hat{\mathbf{G}}_{k,T}=\nabla_{\boldsymbol{\eta}}S_{k,T}(\hat{\boldsymbol{\eta}}_{k}^{(3)})\nabla_{\boldsymbol{\eta}}S_{k,T}(\hat{\boldsymbol{\eta}}_{k}^{(3)})^{\prime}$ and $\hat{\mathbf{D}}_{k,T}=\nabla_{\boldsymbol{\eta}}^{2}S_{k,T}(\hat{\boldsymbol{\eta}}_{k}^{(3)})$ . Hence, a marginal $100\times(1-\alpha)\%$ confidence interval for $\tau_{k}^{0}$ can be constructed by $\hat{\tau}_{k}^{(3)}\pm z_{\alpha/2}\frac{T\sqrt{[\hat{\mathbf{\Sigma}}_{k,T}]_{11}}}{\sqrt{\hat{\tau}_{k+1}^{(3)}-\hat{\tau}_{k-1}^{(3)}}}$, where $z_{\alpha/2}$ is the $(1-\alpha/2)$ quantile of standard normal distribution and $\hat{\mathbf{\Sigma}}_{k,T}]_{11}$ is the element at the last row and last column of the matrix $\hat{\mathbf{\Sigma}}_{k,T}$.